OpenStudy (anonymous):

find the number of three digit numbers that leave remainder 9 when divided by 11

OpenStudy (shane_b):

Basically, your first number will be 108...and then every number +11 after that up to 999: 108, 119, 130, 141, 152, 163, 174, 185, 196, 207, 218, 229, 240, 251, 262, 273, 284, 295, 306, 317, 328, 339, 350, 361, 372, 383, 394, 405, 416, 427, 438, 449, 460, 471, 482, 493, 504, 515, 526, 537, 548, 559, 570, 581, 592, 603, 614, 625, 636, 647, 658, 669, 680, 691, 702, 713, 724, 735, 746, 757, 768, 779, 790, 801, 812, 823, 834, 845, 856, 867, 878, 889, 900, 911, 922, 933, 944, 955, 966, 977, 988

OpenStudy (shane_b):

BTW: That's 81 values in all that meet that criteria.

OpenStudy (anonymous):

1st 3 digit num wch is divisible by 11 is 110 den 121 den 132 n so on.. for remindr 9 u need to add 9 in those numbers.. so 1st 3 digit num wch is divisible by 11 is 110 and last num is 990. 990=110+(n-1)11 ull get value of n=81 if u add 9 to those num,ull get 81... 108 is also a num wch is divisible by 11 nd reminder is 9... so total num's will b 82. and is 82

ganeshie8 (ganeshie8):

yh.. @arch1 is correct. numbers that leave 9 when divided by 11 are in from : 11x + 9 x = [9, 90] satisfies above expression for 3 digit numbers total 82 numbers that match the constraint...

OpenStudy (shane_b):

Yes, I mistakenly left out 999...so it's 82.